RC Meeting, Friday, October 22 , 2021

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 10 am, remote connection link: https://bluejeans.com/861439920

Accelerator Status

  1. Started getting beam back after maintenance day around 00:00 Thursday.
  2. Significant production resumed around 02:00 Thursday.
  3. Beam off about 15:00 on Thursday due to Arc 7 ground fault. Halls C and D down since then.
  4. SL and remote workers for Friday day on standby at the moment.
  5. Pass change for Hall A is now possibly Sunday morning.
  6. And of course the accelerator will be off Monday morning for the Secretary of Energy tour.

Hall C Status

  1. We're waiting for a repair in Arc 7. Anybody got anything they need to do inside Hall C?
  2. So far we have collected about 4127 pions in this configuration

Current Run Plan

  1. Continue data taking until early Nov

Current Issues

  1. Replace flow switches for (HMS?) magnet power supplies: next opportunistic access (Joe Beaufait) ( PENDING )
  2. Q3 insulating vacuum spikes generating alarms. For example, https://logbooks.jlab.org/entry/3930175 .
  3. One HB LHe incident of note: https://logbooks.jlab.org/entry/3929152 . (possible clue to the origin?)
  4. Target ladder camera was identified to be bad: Joe Beaufait will order camera ( PENDING )
  5. Still running with no UPS on the SHMS cryo controls - Steve L. has ordered batteries and UPS. ( PENDING )
  6. HCCAM02 not functioning (lift required to replace).

Special Requests

Useful Links